The chemical analysis of a sample of solid fuel shows it to contain 80.6% carbon, 2.8% hydrogen and 1.1% sulphur. The remainder is incombustible material which remains as ash. Determine (i) the oxygen required for complete combustion of each of the combustible elements per kilogram of fuel and (ii) the theoretical mass of air required for complete combustion per kilogram of the fuel. The oxygen content of air by mass is 23%.
1
Expert's answer
2018-08-15T04:11:09-0400
Question:

w(C)=80.6%=0.806
w(H)=2.8%=0.028
w(S)=1.1%=0.011
m(fuel)= 1 kg = 1000 g
w(O2 in air) = 23%=0.23
m(O2) -?
m(air)-?
________________________________
Solution:
m(C)=m(fuel)×w(C) = 1000×0.806= 806 g
m(H)=m(fuel)×w(H) = 1000×0.028=28 g
m(S)=m(fuel)×w(S) = 1000×0.011=11 g


806g x g
C + O2 = CO2
12 g/mol 32 g/mol
x = (806×32)/12 = 2149 g

28g y g
4H + O2 = 2H2O
4×1g/mol 32g/mol
y = (28×32)/4 = 224 g

11g z g
S + O2 = SO2
32 g/mol 32g/mol
Z = 11×32/32= 11 g
m(O2)= x+y+z = 2149+224+11=2384g
m(air)= m(O2)/w(O2) = 2384/0.23 = 10365 g
Answer: m(O2) = 2384g
m(air)= 10365 g